Sanchez Company has planned capital expenditures that total $2000000. The company wants to maintain a target capital structure that is 35% debt and 65% equity. The company forecasts that its net income this year will be $1,800,000. If the company follows a residual dividend policy, what will be its total dividend payment?
Answer #1

amount financed with equity = 2000000 * 0.65 = 1300000


dividend payment = 1800000 - 1300000 = 500000